OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/4x + 2. y = 4x - 1 y = 4x - 3 y = 1/4x - 3 y = 1/4x - 1

OpenStudy (anonymous):

Use point-slope form, and change it to slope-intercept form to match the form of the answers given.

OpenStudy (anonymous):

Don't make the easy ones hard.\[y+2=\frac{1}{4}(x+4) \implies y=\frac{1}{4}x-1\]